OpenStudy (anonymous):

lim x go 0 cos(x)-1/x =0 who make proof lim x go 0 cos(x)-1/x =0 who make proof @Mathematics

OpenStudy (anonymous):

Well you need to use Squeeze Lemma... in other words. Cos(x) oscillates between 1 and 0. and 1/x approaches zero.

OpenStudy (anonymous):

\[\lim_{x\to0}\frac{\cos(x)-1}{x}\]\[\lim_{x\to0}\frac{-\sin(x)}{1}=0\]

OpenStudy (anonymous):

hımm. thanks for everything

OpenStudy (anonymous):

use Fronklin's i'm sure you are in Calculus 1 and you have not learned l'hopitals which is what mahone is doing